Salt is often used to melt ice and snow on roads in the winter and the temperature right now is - 4.500°C. How much NaCl ( complete dissociation) would be necessary to melt 1000kg of ice by bringing down the freezing temperature to 4.500°C?

1
Expert's answer
2022-03-10T10:46:49-0500

M(NaCl) = 58.5 g/mol;

Kf(H2O) = 1.86 (C * kg/mol)

"\Delta" Tf = b(molality) * Kf;

b = n(solution)/m(solvent) = m(NaCl)/(M(NaCl) * m(H2O));

m(NaCl), kg = b *M(NaCl) * m(H2O)/1000 = "\Delta" Tf *M(NaCl) * m(H2O) / (Kf * 1000)= 4.500 * 58.5 * 1000 / (1.86 *1000) = 141.5 kg.

Answer: 141.5 kg.